OpenStudy (anonymous):

3x -5y = 7 change to y=mx+b

OpenStudy (anonymous):

First notice that in y=mx+b the "x" is on the right side, whereas in 3x-5y=7, the x is on the left...so we have to move it! We can move it over by subtracting 3x from both sides 3x-5y =5 ----> -5y = -3x + 5 Now, you also don't want any numbers in front of the y. Because y is being multiplied here by -5, you should divide through by -5: -5y = -3x + 5 ----> y = 3/5 x - 1